IN 1912, an order by President Taft fixed the width and length of the U.S. flag in the ration 1 to 1.9. If 100 square feet of cloth are to be used to make a U.S. flag, estimate its dimensions to the nearest 1/4 foot.
How can I solve this?

Answer by mananth(16946)

Ratio of width to length = 1:1.9
area of flag = 100 sq. ft
...
Let the constant factor be x
(1*x)*(1.9*x)=100
1.9x^2=100
x^2=100/1.9
x= sqrt(100/1.9)
x= 7.25 feet
...
length = 7.25 *1.9 = 13.77
...
The dimensions are 7.25 feet by 13.75 feet
..
m.ananth@hotmail.ca